# Quantum Reservoir Computing for Realized Volatility Forecasting

### v1

@@ Abstract
Keywords: Realized Volatility Forecasting; Heterogeneous AutoRegressive Models; Quantum Machine Learning; Quantum Reservoir Computing; Explainable AI.
JEL classification: C45, C53, C58
@@ I Introduction
Volatility modeling is a fundamental aspect of financial econometrics, essential to understand and manage the uncertainty or risks associated with financial markets. Accurate modeling and forecasting of volatility are crucial for various applications, including risk management, portfolio optimization, and derivative pricing [1, 2]. Given its critical role, developing robust volatility forecasting models has been a major focus of financial research. While traditional models like Generalized Autoregressive Conditional Heteroskedasticity (GARCH[3] and its variants have been widely adopted, they often struggle to fully capture complex nonlinear dynamics and high-frequency data behaviors characteristic of financial markets [4]. Recognizing these limitations, subsequent research has emphasized the importance of understanding the distribution of realized stock return volatility for more effective financial analysis, as highlighted in Andersen et al. [5]. Furthermore, econometric analysis of realized volatility, particularly its application in estimating stochastic volatility models, has provided crucial insights into the behavior of financial time series, as demonstrated by Barndorff-Nielsen and Shephard [6].
To address some of the shortcomings of traditional models, the Heterogeneous Autoregressive (HAR) model [7] was developed, offering a more nuanced approach by incorporating realized volatility over different time horizons, thus providing a more accurate and comprehensive measure of market risk. Its variations, including HAR-J (HAR with jumps) and CHAR (continuous HAR) [8], SHAR (semivariance-HAR) [9], and HARQ (HAR with realized quarticity) [10], further enhanced its ability to capture different aspects of market volatility.
The nonlinear and often non-Gaussian nature of financial data has driven researchers to explore more sophisticated approaches, such as machine learning techniques to capture complex relationships that traditional econometric models may miss [11, 12, 13, 14, 15, 16, 17, 18, 19]. In the realm of realized volatility forecasting, machine learning models have shown promising results. While some studies found machine learning models to perform similarly to or slightly worse than traditional HAR-family models [20, 21, 22, 23], others reported significant improvements using machine learning approaches [14, 24, 17]. For instance, Christensen et al. [24] applied machine learning to volatility forecasting, demonstrating that these models can significantly improve prediction accuracy, especially when combined with macroeconomic variables. Similarly, Zhang et al. [25] explored the integration of machine learning with intraday commonality, further enhancing the precision of volatility forecasts. [14] also demonstrated the effectiveness of Long Short-Term Memory (LSTM) neural networks in forecasting monthly S&P 500 realized volatility, and have found strong links between volatility and macroeconomic factors paving the way for more advanced machine learning applications in this field. The importance of capturing the multifaceted nature of volatility is underscored by the work of Ghysels et al. [26], who emphasized the value of high-frequency data for more accurate volatility estimation. By leveraging data sampled at different frequencies, their approach offers a richer understanding of market dynamics, which is crucial for effective risk management and portfolio allocation. Recent reviews, such as the one by Gunnarsson et al. [27], have highlighted the growing role of machine learning in volatility modeling, particularly in the prediction of realized and implied volatility indices. These advancements underscore the trend toward more data-driven approaches in finance, which offer superior adaptability to the complexities of modern financial markets. Among machine learning algorithms, recurrent neural networks and their variant, LSTM networks, have shown particular success in modeling time series data due to their inherent ability to capture temporal dependencies. Recurrent neural networks are designed to recognize patterns in sequences of data by maintaining a hidden state that is influenced by previous inputs, making them ideal for time-dependent data such as financial time series [28, 29, 14].
@@
While machine learning has transformed time series modeling, it still operates in the realm of classical computing. On the other hand, the rapidly advancing of quantum computing [30, 31, 32] in recent decades opens up a wholly new frontier. Quantum computers are rapidly emerging in various physical platforms including superconducting qubits [33, 34, 35, 36], ion traps [37, 38, 39, 40], Rydberg atoms [41, 42], photonic setups [43, 44], nitrogen vacancy centers in diamond [45] and topological qubits [46]. While near-term quantum computers have demonstrated quantum supremacy over their classical counterparts, they suffer from various imperfections such as finite coherence time and a limited number of qubits [47]. Therefore, many algorithms with proven quantum advantages cannot be implemented on noisy near-term quantum computers. A natural emergent question is - can near-term quantum computers be used for machine learning? The advent of variational quantum algorithms [48] and related quantum approximate optimization algorithms [49] are attempts to answer this question where they have been adapted to solving problems in a truly diverse array of disciplines, including molecular simulations [50, 51], biology [52], or particle physics [53]. In parallel, methods of classical time series analysis like Crutchfield computational mechanics framework [54, 55], have also been shown to be improved by quantum encodings which reduce model complexity [56]. Quantum computing methods are also starting to become important for modern quantitative finance problems like asset pricing [57], portfolio optimization [58, 59], credit sales classification [60], risk management [61], loan eligibility prediction [62] among others. See Refs  [63, 64, 65] for recent surveys. The key approach behind these algorithms is to optimize a parametrized quantum circuit to optimize the loss function in an iterative way like a classical neural network [48]. However, there is a different paradigm of classical machine learning, namely reservoir computing [66], which does not seek to optimize the parameters of a neural network and training only takes place at the data processing level. Quantum versions of this approach are particularly promising for near-term quantum computers and have been proposed with various underlying platforms like disordered spin chains [67], quantum optical energy levels [68], superconducting devices [69]. They are exceptionally useful for time-series forecasting, as demonstrated with canonical models like Mackey-Glass [70] and autoregressive moving-average (ARMA[71]. Thus, it is natural to ask whether quantum reservoir computing can be useful for accurate forecasting of real-world financial time series data.
@@
This study introduces a novel approach using quantum reservoir computing for realized volatility forecasting of the S&P 500 index. Quantum reservoir computing, an extension of classical reservoir computing, has been chosen for this study due to its ability to efficiently process temporal data [67]. Quantum reservoir computing models harness quantum features to enhance the prediction power of time series data. They utilize the dynamics of fixed quantum systems known as the “reservoir". Unlike conventional neural networks, in reservoir computing, the parameters of the reservoir are not trained, and the learning procedure takes place at the data processing level after performing measurements [72]. This makes reservoir computing particularly advantageous for tasks requiring the capture of temporal dynamics without the heavy computational burden associated with training traditional neural networks. Quantum reservoir computing extends this concept into the quantum domain, leveraging quantum states to enhance computational power and efficiency. This approach can be promising for volatility forecasting, where capturing complex, time-dependent relationships is critical [67, 73, 74, 75, 76, 77, 78]. Such quantum reservoir platforms have already been conceptualized in various atomic and spin lattices [67, 79, 75] as well as photonic platforms [74, 80, 81], It is of particular interest to note that very recent results on reservoir computing already hint that quantum properties may lead to fast and reliable forecasts with smaller resources [82].
Our research aims to showcase the proof of concept and capability of quantum machine learning in real-world time series analysis, particularly in financial forecasting. We benchmark our quantum model against several classical models, including the HAR and HARX models, as well as traditional neural network algorithms. By leveraging a comprehensive set of features, including market microstructure and macroeconomic variables, we explore whether quantum nonlinear models can more effectively capture the intricate relationships governing market volatility. Prior research, such as that by Alaminos et al. [83] and Thakkar et al. [77], has demonstrated the potential of quantum machine learning in financial forecasting, suggesting that quantum models may eventually outperform classical machine learning algorithms, particularly when dealing with large datasets and complex patterns.
@@
The remainder of the paper is organized as follows. Section II outlines the essential technical background, bridging concepts from both financial econometrics and quantum computing to facilitate interdisciplinary understanding. Sec. III demonstrates our method of volatility forecasting with quantum reservoir computing. Sec. IV compares our results with other volatility forecasting strategies, before concluding discussions in Sec. V.
@@ II Technical Background
This paper lies at the intersection of three different subjects, namely econometrics, machine learning, and quantum computation. In this section, we provide a brief overview of these three subjects, introducing the concepts and methodologies that we will use later. Depending on their expertise, the readers can skip the relevant following subsections.
@@ II.1 Classical Models for Realized Volatility Forecasting
where $RV_{t-1}^{(d)}$, $RV_{t-1}^{(w)}$, and $RV_{t-1}^{(m)}$ represent realized volatility averages computed over daily, weekly, and monthly horizons. The coefficients $\beta_{d}$, $\beta_{w}$, and $\beta_{m}$ measure the contributions of these short-term, intermediate-term, and long-term volatility components to the current volatility level. The parameter $\varepsilon_{t}$ is the residual term assumed to follow a conditionally heteroskedastic process, with $\mathbb{E}[\varepsilon_{t}|\mathcal{F}_{t-1}]=0$ and $\operatorname{Var}(\varepsilon_{t}|\mathcal{F}_{t-1})=\sigma_{t}^{2}$, where $\mathcal{F}_{t-1}$ denotes the information set available at time $t-1$. This specification allows the HAR model to capture volatility at multiple frequencies, accommodating heterogeneity in market participants’ investment horizons. The model assumes weak stationarity of the log-volatility series and serially uncorrelated errors. Estimation is typically performed via ordinary least squares on the log-transformed realized volatility series to stabilize variance and reduce the impact of heteroskedasticity. For statistical inference, heteroskedasticity and autocorrelation-consistent (HAC) standard errors are employed following Newey and West [85], ensuring robustness to serial correlation and time-varying error variance.
@@
RVt= β0+ β1RVt-1+ β2(13∑i=13RVt-i)+ β3(112∑i=112RVt-i)+ ∑k=1kγkTXt-k+ εt,
@@
II.2 Quantum Computation Preliminaries
Let us briefly provide an overview of quantum theory necessary for understanding the subsequent sections of this paper. Quantum mechanics offers a mathematical framework originally developed to explain several experimental phenomena that challenged classical Newtonian physics in the late 19th and early 20th centuries. Readers interested in a more comprehensive exploration are encouraged to consult standard references such as Dirac [90] and Nielsen and Chuang [91]. Throughout this paper, we employ the Dirac notation commonly used in quantum physics. To assist readers from financial and econometric backgrounds, we present a minidictionary in Table 1, where the departures of the Dirac notation from the standard econometric notation as laid out in [92] are codified. In modern axiomatic form, quantum mechanics is completely described through the following four axioms.
Table 1: A mini-dictionary of Dirac notation for econometricians.
Object Econometric Notation Dirac Notation Comments
Scalar Variable $a$ $a$
Complex Conjugate of Scalar $\bar{a}$ or $a^{*}$ $a^{*}$
[…]
Adjugate of Matrix A $\textbf{{A}}^{\#}$ $\text{adj}(\mathbf{A})$
Determinant of Matrix A $\det(\textbf{{A}})$ $\det(\mathbf{A})$ Scalar value
[…]
Identity Matrix $\textbf{{I}}_{n}$ $\mathbb{I}$
Zero Vector/Matrix 0 $\mathbf{0}$
[…]
Commutator of Operators $A$ and $B$ $[A,B]=AB-BA$ $[A,B]=AB-BA$ Measures non-commutativity
Anticommutator of Operators $A$ and $B$$\{A,B\}=AB+BA$$\{A,B\}=AB+BA$
Exponential of Operator $\exp(\textbf{{A}})$ $e^{A}$ Defined via power series
[…]
Spin Operators N/A $S_{\alpha}=\frac{\hbar}{2}\alpha$ $\alpha=\{X,Y,Z\}$
Axiom 1: quantum states. – The complete information about a quantum system is encoded in its state vector $|\psi\rangle$, which is a ray of the system Hilbert space $\mathcal{H}$. The quantum bit (or qubit for short) is the smallest nontrivial unit in quantum computing, similar to the bit in classical computing. It is physically implemented by a two-level quantum system, such as the spin of an electron. Its mathematical form can be represented by a state vector in a Hilbert space $\mathcal{H}$, such that the single qubit state is defined as
And,
Axiom 2: composition of quantum systems. – When we consider a joint quantum system consisting of multiple qubits, the state of this system is described by the tensor product Hilbert space of each qubit $\mathcal{H}_{i}$. For example, if two qubits are described by their Hilbert spaces $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ respectively, the joint system comprising both these systems as subsystems is described by the tensor-product Hilbert space $\mathcal{H}_{1}\otimes\mathcal{H}_{2}$. Thus, the state of two qubits is
$|\psi\rangle$ $=|\psi_{1}\rangle\otimes|\psi_{2}\rangle$ (7)
$=\alpha_{00}|00\rangle+\alpha_{01}|01\rangle+\alpha_{10}|10\rangle+\alpha_{11}|11\rangle,$
The dimension of the $n$-qubit system is $2^{n}$, and hence the tensor product space is a $2^{n}$-dimensional complex vector space $\mathcal{C}^{2^{n}}$. The dimension of the $n-$qubit system increases exponentially in the number $n$ of the qubits. Quantum entanglement is another important property of quantum states that is different from the classical states. It occurs when a joint quantum system can not be represented as a tensor product of the subsystems that make up it. Mathematically, if two quantum systems $A$ and $B$ are entangled, the joint system $AB$ might be represented as
$\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|,$ (11)
Axiom 3: quantum dynamics. – The time evolution of a closed quantum system describes how the state of the system evolves over time, which is the most important way to deal with the information. The closed means the evolution of systems that don’t interact with the rest of the world. The time evolution is governed by the Schrödinger equation, which is a fundamental partial differential equation in quantum mechanics.
The time-dependent Schrödinger equation is given by
$i\hbar\frac{d}{dt}|\psi(t)\rangle=H|\psi(t)\rangle,$ (12)
$|\psi(t)\rangle=e^{-\frac{i}{\hbar}Ht}|\psi(0)\rangle,$ (13)
where $|\psi(0)\rangle$ is the initial state of the system at time 00, and $e^{-\frac{i}{\hbar}Ht}$ is the time evolution operator.In practice, it is common to absorb the factor $\hbar$ into $H$, effectively setting $\hbar=1$. If the Hamiltonian $H$ depends on time, the situation is more complex, but we don’t use it in this paper. It is not difficult to see that the time evolution operator $e^{-\frac{i}{\hbar}Ht}$ is a unitary matrix, so this time evolution is a unitary transformation, which preserves the norm of the quantum states, meaning the probability conservation. For simplicity, we usually use $U(t)$ to represent $e^{-iHt}$ as
$|\psi(t)\rangle=U(t)|\psi(0)\rangle.$ (14)
$\rho(t)=U(t)\rho(0)U^{\dagger}(t),$ (15)
Axiom 4: quantum measurement. – Above, we described the evolution of a closed quantum system. However, to gain any information about the system, one has to necessarily consider an experimentalist as an external physical system observing it. Unlike classical mechanics, where in principle, such observations can be made without changing the observed system itself, quantum mechanics distinguishes itself by postulating that the quantum system undergoes irreversible change due to such a measurement. The general quantum measurement postulate is described by a collection measurement operator $\{M_{m}\}$. The index $m$ refers to the measurement outcomes that may occur in the experiment with the probability $p_{m}$ given by the Born rule $p_{m}=\langle\psi|M^{\dagger}_{m}M_{m}|\psi\rangle$, where $|\psi\rangle$ is the state being measured. To ensure the probabilities sum up to unity, the measurement operators satisfy the completeness equation
$\sum_{m}M_{m}^{\dagger}M_{m}=I$ (16)
In most cases of quantum computing, we care about the quantum measurement of an observable. An observable, i.e., a dynamical variable like position or momentum, is described by a self-adjoint operator $A$ (namely $A=A^{\dagger}$), acting on the system Hilbert space $\mathcal{H}$. Via the spectral theorem, any observable has a spectral decomposition
$A=\sum_{m}a_{m}\ket{a_{m}}\bra{a_{m}},$ (17)
where $a_{m}$’s are the eigenvalue and |$\ket{a_{m}}$’s are the eigenvectors of $A$. Note that $\bra{a_{m}}$ | is the dual of the vector $|a_{m}\rangle$, and is an element of the so-called dual space of $\mathcal{H}$ connected to $\mathcal{H}$ via the Riesz representation theorem. For a general quantum state |$\ket{\psi}$, given in Eq. (4), the dual state is given by $\bra{\psi}=[\alpha^{*},\beta^{*}]$, see Table 1. According to quantum physics, the outcome of measuring the operator $A$ on a quantum system described by state |$\ket{\psi}$ probabilistically results in one of the eigenvalues $a_{m}$. The probability of each outcome is determined by the projector $M_{m}{=}\ket{a_{m}}\bra{a_{m}}$ | as $p_{m}=\langle\psi|M^{\dagger}_{m}M_{m}|\psi\rangle=|\langle a_{m}|\psi\rangle|^{2}$. Hence, the expectation value of the measurement is given by
$\langle A\rangle=\sum_{m}p_{m}a_{m}=\langle\psi|A|\psi\rangle$ (18)
@@ II.2 Quantum Computation Preliminaries
II.3 Machine Learning Preliminaries

### v2

@@ I Introduction
Quantum mechanics promises to enhance computing capacity in a fundamental way  [1, 2, 3]. In recent years, quantum computers, albeit with severe limitations, are rapidly emerging in various physical platforms including superconducting qubits [4, 5, 6, 7], ion traps [8, 9, 10, 11], Rydberg atoms [12, 13], photonic setups [14, 15], nitrogen vacancy centers in diamond [16] and topological qubits [17]. While near-term quantum computers have demonstrated quantum supremacy over their classical counterparts, they suffer from various imperfections such as finite coherence time and a limited number of qubits [18]. Therefore, many algorithms with proven quantum advantages cannot be implemented on noisy near-term quantum computers. A natural emergent question is - can near-term quantum computers be used for machine learning? The advent of variational quantum algorithms [19] and related quantum approximate optimization algorithms [20] are attempts to answer this question where they have been adapted to solving problems in a truly diverse array of disciplines, including molecular simulations [21, 22], biology [23], or particle physics [24]. In parallel, methods of classical time series analysis like Crutchfield computational mechanics framework [25, 26], have also been shown to be improved by quantum encodings which reduce model complexity [27]. Quantum computing methods are also starting to become important for modern quantitative finance problems like asset pricing [28], portfolio optimization [29, 30], credit sales classification [31], risk management [32], loan eligibility prediction [33] among others. See Refs  [34, 35, 36] for recent surveys. The key approach behind these algorithms is to optimize a parametrized quantum circuit to optimize the loss function in an iterative way like a classical neural network [19]. However, there is a different paradigm of classical machine learning, namely reservoir computing [37], which does not seek to optimize the parameters of a neural network, and training takes place only at the final output layer.While this approach to machine learning has transformed time series modeling, it still operates in the realm of classical computing [38, 39]. Quantum versions of this approach are particularly promising for near-term quantum computers and have been proposed with various underlying platforms like disordered spin chains [40], quantum optical energy levels [41], spin-boson platforms [42], or superconducting devices [43] for some mathematical and physical problems. They are exceptionally useful for time-series forecasting, as demonstrated with canonical models like Mackey-Glass [44] and autoregressive moving-average (ARMA[45]. Thus, it is natural to ask whether quantum reservoir computing can be useful for accurate forecasting of real-world financial time series data.
Volatility modeling is a fundamental aspect of financial econometrics, essential to understand and manage the uncertainty or risks associated with financial markets. Accurate modeling and forecasting of volatility are crucial for various applications, including risk management, portfolio optimization, and derivative pricing [46, 47]. Given its critical role, developing robust volatility forecasting models has been a major focus of financial research. While traditional models such as Generalized Autoregressive Conditional Heteroskedasticity (GARCH[48] and its standard extensions are widely used, their reliance on low-dimensional parametric recursions for conditional variance, typically linear in squared returns, imposes strong functional restrictions. These constraints may limit their ability to flexibly represent complex nonlinear and multiscale volatility dynamics governing realized volatility at medium and long horizons. Recognizing these limitations, subsequent research has emphasized the importance of understanding the distribution of realized stock return volatility for more effective financial analysis, as highlighted in Andersen et al. [49]. Furthermore, econometric analysis of realized volatility, particularly its application in estimating stochastic volatility models, has provided crucial insights into the behavior of financial time series, as demonstrated by Barndorff-Nielsen and Shephard [50]. To address some of the shortcomings of traditional models, the Heterogeneous Autoregressive (HAR) model [51] was developed, offering a more nuanced approach by incorporating realized volatility over different time horizons, thus providing a more accurate and comprehensive measure of market risk. Its variations, including HAR-J (HAR with jumps) and CHAR (continuous HAR) [52], SHAR (semivariance-HAR) [53], and HARQ (HAR with realized quarticity) [54], further enhanced its ability to capture different aspects of market volatility. The nonlinear and often non-Gaussian nature of financial data has driven researchers to explore more sophisticated approaches, such as machine learning techniques to capture complex relationships that traditional econometric models may miss [55, 56, 57, 58, 59, 60, 61, 62, 63]. In the realm of realized volatility forecasting, machine learning models have shown promising results. While some studies found machine learning models to perform similarly to or slightly worse than traditional HAR-family models [64, 65, 66, 67], others reported significant improvements using machine learning approaches [58, 68, 61]. For instance, Christensen et al. [68] applied machine learning to volatility forecasting, demonstrating that these models can significantly improve prediction accuracy, especially when combined with macroeconomic variables. Similarly, Zhang et al. [69] explored the integration of machine learning with intraday commonality, further enhancing the precision of volatility forecasts. [58] also demonstrated the effectiveness of Long Short-Term Memory (LSTM) neural networks in forecasting monthly S&P 500 realized volatility, and have found strong links between volatility and macroeconomic factors paving the way for more advanced machine learning applications in this field. The importance of capturing the multifaceted nature of volatility is underscored by the work of Ghysels et al. [70], who emphasized the value of high-frequency data for more accurate volatility estimation. By leveraging data sampled at different frequencies, their approach offers a richer understanding of market dynamics, which is crucial for effective risk management and portfolio allocation. Recent reviews, such as the one by Gunnarsson et al. [71], have highlighted the growing role of machine learning in volatility modeling, particularly in the prediction of realized and implied volatility indices. These advancements underscore the trend toward more data-driven approaches in finance, which offer superior adaptability to the complexities of modern financial markets. Among machine learning algorithms, recurrent neural networks and their variant, LSTM networks, have shown particular success in modeling time series data due to their inherent ability to capture temporal dependencies. Recurrent neural networks are designed to recognize patterns in sequences of data by maintaining a hidden state that is influenced by previous inputs, making them ideal for time-dependent data such as financial time series [72, 73, 58].
@@
This work introduces a novel approach using quantum reservoir computing for realized volatility forecasting of the S&P 500 index. Quantum reservoir computing, an extension of classical reservoir computing, has been chosen for this study due to its ability to efficiently process temporal data on near-term quantum computers [40], while maintaining a low training cost and providing a state space that grows exponentially with the number of qubits to capture features. Quantum reservoir computing models harness quantum features to enhance the prediction power of time series data. They utilize the dynamics of fixed quantum systems known as the “reservoir". Unlike conventional neural networks, in reservoir computing, the parameters of the reservoir are not trained, and the learning procedure takes place at the final output layer after performing measurements [74]. This makes reservoir computing particularly advantageous for tasks requiring the capture of temporal dynamics without the heavy computational burden associated with training traditional neural networks. Quantum reservoir computing extends this concept into the quantum domain, leveraging quantum states to enhance computational power and efficiency. This approach can be promising for volatility forecasting, where capturing complex, time-dependent relationships is critical [40, 75, 76, 77, 78, 79, 80]. Such quantum reservoir platforms have already been conceptualized in various atomic and spin lattices [40, 81, 77] as well as photonic platforms [76, 82, 83], It is of particular interest to note that very recent results on reservoir computing already hint that quantum properties may lead to fast and reliable forecasts with smaller resources [84]. We aim to showcase the proof of concept and capability of quantum machine learning in real-world time series analysis, particularly in financial forecasting. We benchmark our quantum model against several classical models, including the HAR and HARX models, as well as traditional neural network algorithms. By leveraging a comprehensive set of features, including market microstructure and macroeconomic variables, we explore whether quantum nonlinear models can more effectively capture the intricate relationships governing market volatility. Prior research, such as that by Alaminos et al. [85] and Thakkar et al. [79], has demonstrated the potential of quantum machine learning in financial forecasting, suggesting that quantum models may eventually outperform classical machine learning algorithms, particularly when dealing with large datasets and complex patterns.
@@
The paper is organized as follows. Section II outlines the classical models, including traditional regression methods and machine learning models, for realized volatility forecasting. Sec. III demonstrates our method of volatility forecasting with quantum reservoir computing. If you are not familiar with quantum computing, Sect. Appendix A gives a brief discussion of the principle of quantum computing, bridging concepts from both financial econometrics and quantum computing to facilitate interdisciplinary understanding. Sec. IV compares our results with other volatility forecasting strategies, before concluding discussions in Sec. V.
@@ II Technical Background
This paper lies at the intersection of three different subjects, namely econometrics, machine learning, and quantum computation. In this section, we provide a brief overview of the former two subjects for physicists, introducing the concepts and methodologies that we will use later. For non-physicists, we include an appendix on quantum computation preliminaries as well as a notation table. Depending on their expertise, the readers can skip the following subsections.
@@ II.1 Classical Models for Realized Volatility Forecasting
where $RV_{t-1}^{(d)}$, $RV_{t-1}^{(w)}$, and $RV_{t-1}^{(m)}$ represent realized volatility averages computed over daily, weekly, and monthly horizons. The coefficients $\beta_{d}$, $\beta_{w}$, and $\beta_{m}$ measure the contributions of these short-term, intermediate-term, and long-term volatility components to the current volatility level. The parameter $\varepsilon_{t}$ is the residual term assumed to follow a conditionally heteroskedastic process, with $\mathbb{E}[\varepsilon_{t}|\mathcal{F}_{t-1}]=0$ and $\operatorname{Var}(\varepsilon_{t}|\mathcal{F}_{t-1})=\sigma_{t}^{2}$, where $\mathcal{F}_{t-1}$ denotes the information set available at time t1t-1. This specification allows the HAR model to capture volatility at multiple frequencies, accommodating heterogeneity in market participants’ investment horizons. The model assumes weak stationarity of the log-volatility series and serially uncorrelated errors. Estimation is typically performed via ordinary least squares on the log-transformed realized volatility series to stabilize variance and reduce the impact of heteroskedasticity. For statistical inference, heteroskedasticity and autocorrelation-consistent (HAC) standard errors are employed following Newey and West [87], ensuring robustness to serial correlation and time-varying error variance.
@@
$\begin{split}RV_{t}&=\beta_{0}+\beta_{1}RV_{t-1}+\beta_{2}\left(\frac{1}{3}\sum_{i=1}^{3}RV_{t-i}\right)+\beta_{3}\left(\frac{1}{12}\sum_{i=1}^{12}RV_{t-i}\right)\\ &\quad+\sum_{k=1}^{k}\gamma_{k}^{T}X_{t-k}+\varepsilon_{t},\end{split}$
@@
II.2 Machine Learning Preliminaries